OpenStudy (anonymous):

Find the partial fraction decompostion -10x/(2x^2-3x-2)

OpenStudy (anonymous):

how to applybthecfunction of sin cos tan in an angle

OpenStudy (anonymous):

\[{-10x \over 2x^2-3x-2}={-10x \over (2x+1)(x-2)}={A \over 2x+1}+{B \over x-2}\] Multiplying both sides by \((2x+1)(x-2)\) gives \(A(x-2)+B(2x+1)=-10x\). Pluging x=2 gives \(5B=-20 \implies B=-4\) and \(x=-1/2\) gives \(-5/2 A=-5 \implies A=2\). So, \(\frac{-10x}{2x^2-3x-2}=\frac{2}{2x+1}-\frac{4}{x-2}\).

OpenStudy (anonymous):

Whoops! A should be -2.

OpenStudy (anonymous):

thanku

OpenStudy (anonymous):

d/dx f[x]

OpenStudy (anonymous):

I don't understand what you're saying santhosh!

OpenStudy (anonymous):

i am not fluent in english sorry

OpenStudy (anonymous):

i want to know more about trignometric calculations

OpenStudy (anonymous):

how to apply the equations

OpenStudy (anonymous):

anwar

OpenStudy (anonymous):

a+b\[\sqrt[?]{?}\]

OpenStudy (anonymous):

OK.. Just write your question in a new post, and I'll help you :)

OpenStudy (anonymous):

\[\sqrt[?]{?}\]